Management is in the process of selecting one out of the following three projects.

Use NPV method to suggest the best option out of the following three projects and give justification for your suggestion. Consider Required Rate of Return as 10%.


Year Cash Flows Project A Project B Project C

0 (Initial Investment) -80 -80 -80

1 (Incomes from Operation Phase of Projects) 30 40 50

2 40 40 40

3 50 40 30


1
Expert's answer
2021-03-29T10:09:48-0400

"NPVa = -80 + 30\/1.1 + 40\/1.1^2 + 50\/1.1^3 = 17.9."

"NPVb = -80 + 40\/1.1 + 40\/1.1^2 + 40\/1.1^3 = 19.47."

"NPVc = -80 + 50\/1.1 + 40\/1.1^2 + 30\/1.1^3 = 21.05."

So, project C is the best option.